# Electron transport in disordered insulating lattice under nonlinear electric field

### v2

@@ IV Conclusion
We investigated the phenomenology of the nonlinear electron transport in a disordered insulator lattice in the steady-state limit under of large electric field. The Keldysh Green’s function method was used to solve the steady-state nonequilibrium rigorously with the dissipation modeled by local fermion thermostats. We set up the lattice in the bulk limit with a large enough self-averaging disorder region. In the experimentally relevant energy scale with the electric field of $10-10^{4}$ kV/cm, we showed that the variable range hopping (VRH) is the dominant transport mechanism when the disorder parametrized by $W$ begins to overcome the gap $\Delta$ in the range $\Delta\lesssim W\lesssim 2\Delta$.
In the VRH regime, the electron drift distance, electric conductivity, and mobility showed strong dependence on $E$ and $W$. In a narrow window of $E$, the hopping range varied more than over an order of magnitude as a function of $E$ and $W$ with a maximum that divides the regime of the VRH and the Bloch oscillation. Conductivity and mobility demonstrated disorder-assisted transport with a huge enhancement over the VRH regime. A comprehensive comparison with the coherent-potential approximation (CPA) highlighted the role of non-local disorder physics in the finite-lattice model.
The nonequilibrium extension of Mott’s VRH argument is confirmed for an electric-field scaling of the conductivity $\sigma(E)$ through an electronic model. Over electric fields outside the linear response and the Bloch oscillation regime, we verified the scaling law $\sigma(E)\sim\exp[-(E_{0}/E)^{\nu}]$ with $\nu=1/3$ and $\nu=1/2$ in the intermediate and strong disorder regime, respectively, in one-dimension. We envision that experimental verification of the nonlinear transport is one of the most pressing issues in nonequilibrium-disorder physics. While the VRH physics has been well documented in the linear regime, its extension to the strong field regime awaits further studies. The transport mechanism in transition-metal oxides, for instance, is often considered as mediated by transport through localized in-gap states, and the nonlinear $IV$ relation before the switching may be examined for the nonlinear transport scaling [54, 46, 49, 48, 20]. Finally, we addressed the VRH mechanism for the unconventional temperature dependence of the conductance threshold which can be experimentally examined in quasi-one-dimensional systems [20, 53].
@@ Acknowledgements.
We thank J. P. Bird, G. Sambandamurthy, M. Randle, H. Zeng, and Z. Zhang for their helpful discussions. In particular we greatly benefited from insightful discussions with B. I. Shklovskii. We acknowledge computational support from the CCR at University at Buffalo. KM and HFF acknowledge the support from the Department of Energy, Office of Science, Basic Energy Sciences, under grant number DE-SC0024139.
@@ Appendix A Nonlinear Transport in Clean Limit
Adopting the similar approach in the previous section, we can write the local electron density $n_{\rm loc}$ in the electric field in terms of the same-time lesser Green’s function, after the change of variable $p+ES\to p$ in the expression defined in Appendix A of Paper I, as
@@
[…]
@@
[…]
@@
[…]
@@
[…]
×$\times\exp\left[is\epsilon(p)+\frac{iE^{2}s^{3}}{24m}\right]$
@@
×$\times\int_{-\infty}^{\infty}ds\frac{1+\gamma|s|}{(s+i\eta)^{3/2}}e^{-\gamma|s|+i\lambda(s+s^{3}/3)}.$
@@
[…]

### v3

@@ IV Conclusion
We investigated the phenomenology of the nonlinear electron transport in a disordered insulator lattice in the steady-state limit under a large electric field. The Keldysh Green’s function method was used to solve the steady-state nonequilibrium rigorously with the dissipation modeled by local fermion thermostats. We set up the lattice in the bulk limit with a large enough self-averaging disorder region. In the experimentally relevant energy scale with the electric field of $10-10^{4}$ kV/cm, we showed that the variable range hopping (VRH) is the dominant transport mechanism when the disorder parametrized by $W$ begins to overcome the gap $\Delta$ in the range $\Delta\lesssim W\lesssim 2\Delta$.
In the VRH regime, the electron drift distance, electric conductivity, and mobility showed strong dependence on $E$ and $W$. In a narrow window of $E$, the hopping range varied more than an order of magnitude as a function of $E$ and $W$, with a maximum that divides the regime of the VRH and the Bloch oscillation. Conductivity and mobility demonstrated disorder-assisted transport with a huge enhancement over the VRH regime. A comprehensive comparison with the coherent-potential approximation (CPA) emphasized the role of non-local disorder physics in the disordered-lattice model.
The nonequilibrium extension of Mott’s VRH scenario is confirmed by an electric-field scaling of the conductivity $\sigma(E)$ through an electronic model. For electric fields, outside the linear response and the Bloch oscillation regime, we verified the scaling law $\sigma(E)\sim\exp[-(E_{0}/E)^{\nu}]$ with $\nu=1/3$ and $\nu=1/2$ in the intermediate and strong disorder regime, respectively, in one-dimension. This can be attributed to electric field-driven electrons occupying and hopping over non-local localized sites in a similar energy range in the disordered active region. We envision that experimental verification of the nonlinear transport is one of the most pressing issues in nonequilibrium-disorder physics. While the VRH physics has been well documented in the linear regime, its extension to the strong field regime awaits further studies. The transport mechanism in transition-metal oxides, for instance, is often considered as mediated by transport through localized in-gap states, and the nonlinear $IV$ relation before the switching may be examined for the nonlinear transport scaling [55, 49, 45, 46, 20]. Finally, we addressed the VRH mechanism for the unconventional temperature dependence of the conductance threshold, which can be experimentally examined in quasi-one-dimensional systems [20, 54].
@@ Acknowledgements.
We thank J. P. Bird, J. Hofmann, G. Sambandamurthy, M. Randle, H. Zeng, and Z. Zhang for their helpful discussions. In particular we greatly benefited from insightful discussions with B. I. Shklovskii. We acknowledge computational support from the CCR at University at Buffalo. KM and HFF acknowledge the support from the Department of Energy, Office of Science, Basic Energy Sciences, under grant number DE-SC0024139.
@@ Appendix A Nonlinear Transport in Clean Limit
Adopting a similar approach to our previous paper [34], we can write the local electron density $n_{\rm loc}$ in the electric field in terms of the same-time lesser Green’s function, after the change of variable $p+ES\to p$ in the expression defined in Appendix A [Eq.(A6) of [34]], as
@@
$=$ $\frac{i}{4\pi^{2}}\int dp\int_{-\infty}^{\infty}ds\frac{e^{-\Gamma|s|}}{s+i\eta}\exp\left[is\epsilon(p)+\frac{iE^{2}s^{3}}{24m}\right].$
@@
$\frac{ie^{i\pi/4}m^{1/2}}{(2\pi)^{3/2}}\int_{-\infty}^{\infty}ds\frac{e^{-\Gamma|s|}}{(s+i\eta)^{3/2}}\exp\left[i\Delta s+\frac{iE^{2}s^{3}}{24m}\right].$ (28)
@@
$ie^{i\pi/4}\left(\frac{\sqrt{m}E}{(2\pi)^{3}\sqrt{8\Delta}}\right)^{1/2}\int_{-\infty}^{\infty}ds\frac{e^{-\gamma|s|}}{(s+i\eta)^{3/2}}e^{i\lambda(s+s^{3}/3)},$ (29)
@@
$J(E)$ $=$ $\frac{iE}{8\pi^{2}m\Gamma}\int dp\int_{-\infty}^{\infty}ds\frac{(1+\Gamma|s|)e^{-\Gamma|s|}}{s+i\eta}$
$\times\exp\left[is\epsilon(p)+\frac{iE^{2}s^{3}}{24m}\right]$
@@
$\times\int_{-\infty}^{\infty}ds\frac{1+\gamma|s|}{(s+i\eta)^{3/2}}e^{-\gamma|s|+i\lambda(s+s^{3}/3)}.$
@@
$\int_{-\infty}^{\infty}ds\frac{1-\frac{1}{2}\gamma^{2}s^{2}}{(s+i\eta)^{3/2}}e^{i\lambda(s+s^{3}/3)}$ (32)